# Investigating learning-independent abstract reasoning in artificial neural networks

### v1

@@ Abstract
Humans are capable of solving complex abstract reasoning tests. Whether this ability reflects a learning-independent inference mechanism applicable to any novel unlearned problem or whether it is a manifestation of extensive training throughout life is an open question. Addressing this question in humans is challenging because it is impossible to control their prior training. However, assuming a similarity between the cognitive processing of Artificial Neural Networks (ANNs) and humans, the extent to which training is required for ANNs’ abstract reasoning is informative about this question in humans. Previous studies demonstrated that ANNs can solve abstract reasoning tests. However, this success required extensive training. In this study, we examined the learning-independent abstract reasoning of ANNs. Specifically, we evaluated their performance without any pretraining, with the ANNs’ weights being randomly-initialized, and only change in the process of problem solving. We found that naive ANN models can solve non-trivial visual reasoning tests, similar to those used to evaluate human learning-independent reasoning. We further studied the mechanisms that support this ability. Our results suggest the possibility of learning-independent abstract reasoning that does not require extensive training.
@@ Introduction
Recent developments in the field of machine learning are revolutionizing the cognitive sciences because, for the first time, we have access to non-human agents capable of performing complex cognitive functions [1]. These machines are constructed of deep Artificial Neural Networks (ANNs) trained using stochastic gradient descent methods on large datasets [2]. Whether or not these machines exhibit “intelligence” in a similar way to humans’ intelligence is still a matter of debate [3, 4, 5]. What is clear, however, is that the training of humans and machines differs on several levels. One of the prominent differences is the size of the training set. For example, current Large Language Models (LLM) pretrain on datasets that are orders of magnitudes larger than those children are exposed to [6], a difference that echoes Chomsky’s arguments several decades ago about “poverty of the stimulus” in language learning [7].
@@
The topic of this paper is abstract reasoning, sometimes referred to as “fluid intelligence” [8]. Abstract reasoning is, broadly speaking, the ability to solve complex problems by identifying regularities and relations in the problem being solved and utilizing them for deducing the solution [9, 10]. It is often studied using intelligence tests that comprise word analogy tests (e.g., infer that the relationship between “cow” and “milk” is the same as between “chicken” and “egg”) and visual reasoning tests (e.g., Raven Progression Matrices) [11, 12]. The extent to which abstract reasoning in humans relies on prior training is a matter of debate [13]. In support of a capacity for learning-independent abstract reasoning is the observation that individuals’ performances on different tests are correlated [14]. Because different tests seem to require different prior knowledge, these correlations are taken as support for the hypothesis that a single learning-independent ability, “abstract reasoning”, is used for all these various tasks. Moreover, this general “abstract-reasoning” ability does not seem to improve by practice – training on one task does not generalize to other tasks [15].
Strong support for the hypothesis that abstract reasoning requires extensive training comes from machine learning. Recent developments in this field demonstrate that extensively-trained deep networks can successfully solve many types of intelligence tests [16, 17, 18, 19, 20, 21]. The performance of these networks relies on extensive prior training, suggesting that their success in solving abstract reasoning benchmarks might be related to exposure to similar questions (an issue called contamination) [3, 4], thus memory-dependent. Indeed, changes in how problems are phrased that seem minor to humans and do not affect their performance can render them unsolvable to the networks [5].
Trivially, solving some forms of intelligence tests requires some form of training. A human who does not speak English or has never heard about the relationship between “cow” and “milk” will not be able to analogously relate “chicken” to “egg”. The extent to which such training is required for solving visual reasoning tests is less clear. Visual reasoning tests utilize abstract shapes like squares and triangles that are presumably more general than language. However, even this is unclear because the everyday visual experience may be a form of extensive training [22].
@@
To contribute to this debate, here we study the extent to which ANN models can solve visual abstract reasoning tests without prior training. To that goal, we constructed visual abstract reasoning tests that require identifying relations in a sequence of stimuli, a sub-task shared among intelligence tests [9, 10, 8, 23]. We studied the ability of Relation Networks (RNs) [24] to solve them, as members of this class of models were shown to be capable of identifying abstract relations and solving intelligence tests after extensive training [25, 18]. Furthermore, RNs were shown to successfully solve word-analogy problems without specific training on those problems, relying on general pretrained relations [26, 27]. In contrast to these previous studies, our focus was the ability of “naive” RNs, who were not exposed to any pre-training, to identify relations in visual reasoning tests and use them for solving the tests.
The paper is constructed as follows: We introduce the problems and the model that we tested. Then, we present its performance and dissect the underlying requirements for the model’s successful problem-solving. Finally, we add training to the model and discuss a similarity between the effect of the training schedule on performance and a phenomenon known as the interleaving effect in the cognitive sciences.
@@ Results
We constructed a set of artificial problems in which the task is to evaluate the consistency of an input with a sequence of its preceding inputs (Fig. 1). Each problem comprises $5$ gray-scale images and $4$ optional-choice images. The images, $224\times 224$ pixels each, are composed of identical abstract objects and differ along several dimensions: the shape of the objects, their size, their color, their number, and their arrangement. By construction, one of these features changes predictably over the $5$ images. Formally, an image is characterized by a low-dimensional vector of features, $\mathbf{f}_{j}$, where $f^{i}_{j}$ denotes the value of feature $i$ in image $j$. An image $\mathbf{x}_{j}$ is constructed according to its characterizing features by a generative function $\mathbf{x}_{j}=\mathbf{G}\left(\mathbf{f}_{j}\right)$. One of the features $f^{p}$ changes predictably along the sequence according to a simple deterministic rule $f_{j+1}^{p}=U(f_{j}^{p})$ while the other features are either constant over the images or change randomly (values are i.i.d). Considering the optional-choice images, the predictable rule is followed in only one of them, and the task is to select this image. The other features are either constant in all $9$ images ($5$ of sequence and $4$ of optional choices) or change randomly (see Methods). We refer to a problem’s predictably changing feature as the problem’s Predictive Feature (PF) and to the randomly changing features as distractors. Intuitively, the number of distractors is a measure of a problem’s difficulty.
@@
This loss is minimized for a relation function $R_{\theta}\left(Z_{\phi}\left(\mathbf{x}_{i}\right),Z_{\phi}\left(\mathbf{x}_{j}\right)\right)$ that outputs a minimal relation score for consecutive sequence images ($j=i+1$), requiring the identification of the regularity that characterizes these consecutive images. We updated the networks’ weights $\theta$ and $\phi$ with 10 optimization steps over the loss $\mathcal{L}\left(\theta,\phi\right)$ using the RMSprop optimizer [28] (learning rate of $1\mathrm{e}{-5}$, the rest of the parameters are set to PyTorch [29] default). Eventually, after optimization, we evaluated the consistency of each choice image with the sequence based on their relation value $R$ when they were placed as the sixth sequence image $R_{\theta}\left(Z_{\phi}\left(\mathbf{x}_{5}\right),Z_{\phi}\left(\ \cdot\ \right)\right)$ and selected the choice image with the lowest relation value as the answer.
@@
We evaluated the performance of the vanilla model on the different tests, in which the predictive feature’s values increased linearly, and found that it performed substantially better than chance ($0.25$) in almost all tasks and all levels of difficulty (Fig. 2). Without distractors, its performance on some tasks was close to perfect. We also found that performance decreased with the number of distractors, verifying that the number of distractors is a good measure of the task’s difficulty. All these results were obtained using networks that we randomly initialized before each problem, thus demonstrating that learning-independent abstract reasoning is possible, at least to some extent. Averages over all conditions, the model’s performance was $0.58\pm 0.01$. From this point in the paper, we use this global performance measure for comparisons (see Methods; complete performance results are in the Supplementary Information).
@@
Together, our results indicate that the ANN’s ability to execute abstract reasoning without learning stems from two important properties: (1) The random convolutional layers extract features that are correlated with the relevant features. (2) The optimization process amplifies the response to those features that monotonically vary along the sequence.
@@
To conclude this section, we demonstrated two ways for carrying the computational load: either the encoder carries most of the load by extracting the relevant feature such that a simple linear relation module is sufficient for capturing multiple monotonic rules. Alternatively, the relation module can carry the computational load. In that case, the relation module finds a specific relation between high-dimensional input representations, keeping the input representations fixed during problem-solving and usable for finding additional relations.
@@ Discussion
The main result of this paper is that abstract reasoning is possible even in naive (and random) ANNs. This is a result that has implications both for the cognitive sciences and for machine learning. In the cognitive sciences, whether abstract reasoning is a learning-independent capability or whether it requires extensive training has long been a matter of contention [13, 7]. It is difficult to address this question in humans, because our interaction with the environment, both in development and in evolutionary time scales, is a form of extensive training. Our work shows that some non-trivial aspects of abstract reasoning are possible (at least in ANNs) without prior learning. In machine learning, much progress in recent years was based on using ever larger datasets to train ever more complicated ANNs [35, 36]. Our work shows that ANNs are computationally powerful even in the absence of large datasets.
Traditionally, abstract reasoning in the brain was considered a symbolic computation: a digital computation that is very different from the analog computation that ANNs perform [37, 38]. More recently, however, it was demonstrated that computations that were attributed to complex symbolic manipulations could be achieved by extensively-trained ANNs [39, 40]. This ability is particularly pronounced when considering LLMs, that seem capable of abstract reasoning [21]. Criticizing abstract reasoning in ANNs, it has been argued that the success of these networks reflects effective memory-retrieval rather than ”true“ abstract reasoning [3, 4]. Our contribution is that we show that such networks can achieve what looks like symbolic abstract reasoning without any training and, hence, without any memory recall.
Fundamental to the ability of our network to perform the task is the encoder, which extracts features that co-vary with the images. The random convolutional layers were found to be instrumental in extracting features that are correlated with relevant latent features. Highlighting the features that are relevant for the particular sequence of images of a particular problem is done in the deep layers of the encoder, together with the relation module.
@@
These results resonate with human-brain studies. In humans, the primary regions in the visual cortex are considered generic feature extractors that are sensitive to low-level features of the stimulus, such as orientation and direction. It seems unlikely that this low-level feature extraction changes with every problem presented to a human participant (they may, however, change with extensive training [41]). Imaging studies have revealed that higher cortical regions, such as the lateral prefrontal cortex, play an important role in abstract reasoning [42] and rule learning [43]. We hypothesize that these higher cortical regions perform the computation of highlighting the relevant feature (Fully-connected layers of the encoder) and identifying its regularity (the relation module).
@@
Notably, the computations performed by the fully-connected layers of the encoder and the relation module are somewhat interchangeable. This is because either of those networks could carry the main computational load. This suggests an interesting approach for finding relations by implementing a few very simple and general (applicable to different problems) relation modules, transferring a significant computational load of finding appropriate input representations to the encoder. Given the interchangeability of complexity in the fully-connected layers of the encoder and the relational module, the separation between the encoder and relational module in the brain may be somewhat artificial. A clear border between an encoder and a relational module may not exist.
@@
Our framework naturally generalizes to explaining knowledge acquisition through problem-solving (in an unsupervised way, without feedback). We found that training with many short interleaved blocks was substantially more effective than training with two long blocks. This resembles a similar observation in the cognitive sciences known as the interleaving effect [33, 34]. In the cognitive sciences, two competing theories have been used to explain this effect. In one, the interleaving effect is due to enhanced problem identification and feature distinction required when solving two types of problems in close proximity [44]. The second theory explains the interleaving effect by proposing that with interleaved training, the brain is continually engaged at retrieving the responses from memory – a process that enhances the consolidation of those memories [45, 46]. In contrast to these theories, our model has no explicit problem-identification or memory-consolidation mechanisms implemented. Rather, it is another manifestation of the well-known catastrophic forgetting phenomenon in machine learning [31, 32].
@@
While explaining some aspects of abstract reasoning, our model does not encapsulate other facets of abstract reasoning observed in humans. Specifically, our model does not incorporate working memory, limiting the regularities it can identify. It also does not explicitly perform the mapping computation required for analogical reasoning [27]. Additionally, the model cannot solve a problem by breaking it into its sub-components [47]. For example, to solve a Raven Progression Matrix, humans use the strategy of identifying common regularities in the rows and the columns. Our model was constructed only to find a regularity in a sequence. As with most ANN models, the model cannot interpret its choices. Finally, it lacks the ability to generate new images that follow the regularity it identifies. These limitations present opportunities for future research and suggest areas for improvement.
Regarding human abstract reasoning, the recent success of extensively trained deep learning models in solving abstract reasoning problems supports the hypothesis that human abstract reasoning requires extensive training. Our work shows that similar deep learning models can also exhibit some form of learning-independent abstract reasoning. As deep networks are currently considered leading models for human cognition, our work indicates that learning-independent abstract reasoning is also possible in humans.

### v2

# Untrained neural networks can demonstrate memorization-independent abstract reasoning

@@ Abstract
The nature of abstract reasoning is a matter of debate. Modern artificial neural network (ANN) models, like large language models, demonstrate impressive success when tested on abstract reasoning problems. However, it has been argued that their success reflects some form of memorization of similar problems (data contamination) rather than a general-purpose abstract reasoning capability. This concern is supported by evidence of brittleness, and the requirement of extensive training. In our study, we explored whether abstract reasoning can be achieved using the toolbox of ANNs, without prior training. Specifically, we studied an ANN model in which the weights of a naive network are optimized during the solution of the problem, using the problem data itself, rather than any prior knowledge. We tested this modeling approach on visual reasoning problems and found that it performs relatively well. Crucially, this success does not rely on memorization of similar problems. We further suggest an explanation of how it works. Finally, as problem solving is performed by changing the ANN weights, we explored the connection between problem solving and the accumulation of knowledge in the ANNs.
@@ Introduction
The topic of this paper is abstract reasoning, sometimes referred to as “fluid intelligence” [1]. Abstract reasoning is, broadly speaking, the ability to solve complex problems by identifying regularities and relations in the problem being solved and utilizing them for deducing the [2, 3]. It is often studied using intelligence tests that comprise word analogy tests (e.g., infer that the relationship between “cow” and “milk” is the same as between “chicken” and “egg”) and visual reasoning tests (e.g., Raven Progression Matrices) [4, 5]. As artificial intelligence continues to advance, understanding the nature of abstract reasoning in both humans and machines is becoming a central question in cognitive science and AI research [6].
Abstract reasoning in artificial neural networks (ANNs) appears to be closely tied to training. While deep ANNs have shown impressive performance on various intelligence tests [7, 8, 9, 10, 11, 12], their success relied heavily on extensive prior training. Additionally, questions have been raised about the nature of this performance. There are indications that ANNs’ success may stem more from “contamination” – exposure to similar questions in their training data – rather than from genuine abstract reasoning [13, 14]. This dependency on specific training data is further emphasized by findings that minor changes in problem phrasing, which do not affect human performance, can render problems unsolvable for ANNs [15, 6]. Thus, while ANNs may exhibit some analogical reasoning capabilities, it is disputed that these are based on pattern matching or memorization rather than on general intelligence comparable to that of humans.
In this work, we investigated whether ANN tools commonly used in machine learning are capable of demonstrating general abstract reasoning. Specifically, we asked if these networks could solve intelligence test problems with novel inputs, relying only on the information provided by the specific problem at hand, without drawing on prior memorization.
Certain intelligence tests, by their nature, require some level of prior knowledge. For instance, a human unfamiliar with English or the relationship between “cow” and “milk” would struggle to relate “chicken” to “egg” in an analogy test. Consequently, general intelligence in humans is often assessed using visual reasoning tests, utilizing abstract shapes like squares and triangles to minimize the influence of language or cultural knowledge. Thus, to evaluate the abstract reasoning of ANN models, we employed visual abstract reasoning tests. These visual reasoning tests require identifying relations in a sequence of stimuli, a skill common to many intelligence tests [2, 3, 1, 16].
@@
For our network models, we used Relation Networks (RNs) [17], as members of this class of models were shown to be capable of identifying abstract relations and solving intelligence tests after extensive training [18, 9]. Notably, RNs were shown to successfully solve word-analogy problems without specific training on those problems, but with specific training on relevant relationships [19, 20]. In contrast to these previous studies, our focus was on the ability of “naive” RNs, who were not exposed to any pre-training, to identify relations in visual reasoning tests and use them for solving the tests.
The structure of this paper is as follows: We begin by introducing the visual reasoning tests and the network models employed in our study. Next, we present the model’s performance, showcasing its ability to solve non-trivial problems. We then analyze the mechanisms underlying this performance. Finally, as the model’s problem-solving involves an optimization process that modifies network parameters in a manner similar to learning, we examine the relationship between the model’s problem-solving capabilities and the networks’ accumulation of knowledge.
@@ Results
We constructed a set of artificial problems in which the task is to evaluate the consistency of an image with a sequence of its preceding images (Fig. 1). Each problem comprises $5$ gray-scale images and $4$ optional-choice images. The images, $224\times 224$ pixels each, are composed of identical abstract objects and differ along several dimensions: the shape of the objects, their size, their color, their number, and their arrangement. By construction, one of these features changes predictably over the $5$ images. Formally, an image is characterized by a low-dimensional vector of features, $\mathbf{f}_{j}$, where $f^{i}_{j}$ denotes the value of feature $i$ in image $j$. An image in pixel space, $\mathbf{x}_{j}$, is constructed according to its characterizing features by a generative function $\mathbf{x}_{j}=\mathbf{G}\left(\mathbf{f}_{j}\right)$. One of the features $f^{p}$ changes predictably along the sequence according to a simple deterministic rule $f_{j+1}^{p}=U(f_{j}^{p})$ while the other features are either constant over the images or change randomly (values are i.i.d). Considering the optional-choice images, the predictable rule is followed in only one of them, and the task is to select this image. The other features are either constant in all $9$ images ($5$ of sequence and $4$ of optional choices) or change randomly (see Methods). We refer to a problem’s predictably changing feature as the problem’s Predictive Feature (PF) and to the randomly changing features as distracting features or distractors. Intuitively, the number of distractors is a measure of a problem’s difficulty.
@@
This loss is minimized for a relation function $R_{\theta}\left(Z_{\phi}\left(\mathbf{x}_{i}\right),Z_{\phi}\left(\mathbf{x}_{j}\right)\right)$ that outputs a minimal relation score for consecutive sequence images ($j=i+1$), requiring the identification of the regularity that characterizes these consecutive images. We updated the networks’ weights $\theta$ and $\phi$ with 10 optimization steps over the loss $\mathcal{L}\left(\theta,\phi\right)$ using the RMSprop optimizer [21] (learning rate of $10^{-5}$, the rest of the parameters are set to PyTorch [22] default). Eventually, after optimization, we evaluated the consistency of each choice image with the sequence based on their relation value $R$ when they were placed as the sixth sequence image $R_{\theta}\left(Z_{\phi}\left(\mathbf{x}_{5}\right),Z_{\phi}\left(\ \cdot\ \right)\right)$ and selected the choice image with the lowest relation value as the answer.
@@
We evaluated the performance of the vanilla model on the different tests, in which the predictive feature’s values increased linearly, and found that it performed substantially better than chance ($0.25$) in almost all tasks and all levels of difficulty (Fig. 2). Without distractors, its performance on some tasks was close to perfect. We also found that performance decreased with the number of distractors, verifying that the number of distractors is a good measure of the task’s difficulty. All these results were obtained using networks that we randomly initialized before each problem, thus demonstrating that ANNs can perform abstract reasoning that does not depend on memorization. Averages over all conditions, the model’s performance was $\mathbf{0.58\pm 0.01}$. From this point in the paper, we use this global performance measure for comparisons (see Methods; complete performance results are in the Supplementary Information).
@@
Together, our results indicate that the ANN’s ability to execute abstract reasoning without prior learning stems from two important properties: (1) The random convolutional layers extract features that are correlated with the relevant features. (2) The optimization process amplifies the response to those features that monotonically vary along the sequence.
@@
To conclude this section, we demonstrated two ways for carrying the computational load. Either the encoder carries most of the load by extracting the relevant feature in a manner that a simple linear relation module is sufficient for capturing the rule. Alternatively, the relation module can carry the computational load. In that case, the relation module finds a specific relation between high-dimensional input representations, keeping the input representations fixed during problem-solving.
@@ Discussion
We found that naive randomly-initialized ANNs can perform abstract reasoning that does not rely on memorization when they are optimized at test time. This result has implications both the cognitive sciences and for machine learning.
Traditionally, abstract reasoning in humans has been considered a symbolic computation – a type of digital processing distinctly different from the analog nature of computation in ANNs [29, 30]. Recently, however, studies have shown that complex computations once attributed solely to symbolic processing can be accomplished by extensively trained ANNs [31, 32]. This is especially evident with large language models, which appear capable of performing certain forms of abstract reasoning [12]. Nevertheless, critics of abstract reasoning in ANNs argue that this success may be due more to sophisticated memory retrieval than to genuine abstract reasoning [13, 14]. Our contribution is that we show that the tools used for training ANNs can also be used for exhibiting what resembles symbolic abstract reasoning without any training, hence without relying on memory recall.
@@
A key element in our network’s ability to perform abstract reasoning tasks is the convolutional part of the encoder. We found that these random convolutional layers are instrumental in extracting features correlated with relevant latent features. Interestingly, optimization of the convolutional layers was not necessary for achieving the performance. These results resonate with human-brain studies. In humans, early visual cortex regions act as general-purpose feature extractors, sensitive to basic features like orientation and direction. It seems unlikely that this low-level feature extraction changes with every problem presented to a human participant. They may, however, change with extensive training [33].
Highlighting the features that are relevant for the particular sequence of images of a particular problem was done in the deep layers of the encoder (fully-connected layers), together with the relation module. In humans, imaging studies suggest that higher cortical regions, such as the lateral prefrontal cortex, play an important role in abstract reasoning [34] and rule learning [35]. We hypothesize that these higher cortical regions perform the analog of the optimization-based computation of highlighting the relevant feature (fully-connected layers of the encoder) and identifying its regularity (the relation module).
@@
Notably, the computations performed by the fully-connected layers of the encoder and the relation module are somewhat interchangeable. This is because either of those networks could carry the main computational load. This suggests an interesting approach for finding relations by implementing a few very simple and general (applicable to different problems) relation modules, transferring a significant computational load of finding appropriate input representations to the encoder. Furthermore, given the interchangeability of complexity in the fully-connected layers of the encoder and the relational module, the separation between the encoder and relational module in the brain analog of this computation may not exist.
@@
Our framework naturally generalizes to explaining knowledge acquisition through problem-solving (the practice in Fig. 7 was unsupervised, with no feedback). We found that training with many short interleaved blocks was substantially more effective than training with two long blocks. This resembles a similar observation in the cognitive sciences known as the interleaving effect [27, 28]. In the cognitive sciences, two competing theories have been used to explain this effect. In one, the interleaving effect is due to the enhanced problem-identification and feature-distinction required when solving two types of problems in close proximity [36]. The second theory explains the interleaving effect by proposing that with interleaved training, the brain is continually engaged at retrieving the responses from memory – a process that enhances the consolidation of those memories [37, 38]. In contrast to these theories, our model has no explicit problem-identification or memory-consolidation mechanisms implemented. Rather, the interleaving effect is a manifestation of the well-known catastrophic forgetting phenomenon in machine learning [25, 26].
We focused in this paper on the abstract reasoning of ANNs optimized by gradient descent. These models have shown to achieve performance levels that sometimes rival or even exceed human capabilities, especially in areas like language and visual processing, key aspects of human cognition [10, 39]. However, these successes have been achieved by scaling up model size and training data, with models trained on datasets vastly larger than what human children require to learn comparable skills [40, 41]. Thus, the sustainability of simply increasing network size and data volume as a path to further improvements has been doubted [42, 43]. Our findings suggest a potential alternative approach, in which ANNs, by optimizing their weights at test time, exhibit computational capacity in the absence of massive datasets. This approach offers a promising direction for AI development that prioritizes efficiency over scale.
@@
Abstract reasoning consists of several computational facets. In this work, we focused on only one of them: the identification of relationships between images in order to infer the sequence completion, also termed inductive reasoning. Inductive reasoning is needed for solving many types of intelligence tests [16]. While modeling this facet, our model does not encapsulate other facets of abstract reasoning observed in humans. Specifically, our model does not incorporate working memory, limiting the regularities it can identify. It also does not explicitly perform the mapping computation required for analogical reasoning [20]. Additionally, the model cannot solve a problem by breaking it into its sub-components [44]. For example, to solve a Raven Progression Matrix, humans use the strategy of identifying common regularities in the rows and the columns. Our model was constructed only to find a regularity in a sequence. As with most ANN models, the model cannot interpret its choices. Finally, it lacks the ability to generate new images that follow the regularity it identifies. These limitations present opportunities for future research and suggest areas for improvement.
In humans, evidence suggests that abstract reasoning operates as a general computational process, analogous to a general-purpose computer that can handle any input. For example, an individual’s performance on various cognitively demanding tests tends to correlate [45]. As the tests require different prior knowledge, these correlations are taken as support for the hypothesis that a general ability, often termed general intelligence [46], underlies these diverse cognitive skills. Additionally, training on a specific cognitive task usually does not improve performance on unrelated tasks [24]. This lack of transfer suggests that human abstract reasoning is indeed general, relying on general cognitive processes rather than specific learned patterns or memorized solutions. This somewhat resembles our model.
In conclusion, our work demonstrates that ANNs can exhibit abstract reasoning abilities without reliance on memory recall, opening pathways for further exploration of abstract reasoning mechanisms in both artificial systems and humans.